A contractor has purchased a wheeled loader for $130,000 and expects to use the loader an average of 1,500 hours per year. Tires cost $6,000 to replace (estimated to occur after each 4,500 hours of use), and major repairs will be needed every 6,000 hours at a cost of $5,000. The contractor expects to be able to sell the loader for $10,000 after she has used it for 15,000 operating hours. Fuel, oil, and minor maintenance cost about $19.75 for each hour the loader is used. Interest, insurance, and taxes total about 16%. How much should the contractor charge per hour for use of the loader to recover her costs?

How to do it on the TMV solver on a Ti-84
Answer #1

As per Chegg Guidelines we cannot provide direct answers. However, I've answered it as below
Violating them will get my account suspended.
So please understand and like this answer

Wheeled loader cost for $130,000

Average of 1,500 hours per year.

Tires cost $6,000 to replace (estimated to occur after each 4,500 hours of use), and major repairs will be needed every 6,000 hours at a cost of $5,000. The contractor expects to be able to sell the loader for
$10,000 after she has used it for 15,000 operating hours. Fuel, oil, and minor maintenance
cost about $19.75 for each hour the loader is used. Interest, insurance, and
taxes total about 16%.

Selling Price= $ 10,000

Time=10 yrs =15000 operating hours

Total cost for 15000 operating hour=130,000+tire cost+maintenance+19.75*15000+0.16*130,000=130000+18000+10000+296250+20800 =$ 475050

So total cost to recover=475050-10000=465050

Average charge per hour=465050/15000=$ 31.03